# code-for-first-row = , code-for-last-row = , code-for-first-col = , code-for-last-col =

### v1

@@ 8.3 Neurocognition Modelling
Finite partially ordered sets are natural models for cognition, as it is reasonable to assume that some cognitive states have higher levels of functionality than others[165, 166, 164, 169, 170, 170]. Finite partially ordered classification models are useful for many statistical applications, including cognitive modelling. A data analytic framework for implementing latent finite partially ordered classification models is introduced which proposes such tools which is useful cognitive applications that are latent and complex. Poset models are flexible and can become quite rich and complex, enabling them to be effective models for describing response phenomena from test educational test data or neuropsychological assessment data[165, 167, 173]. An objective of neuropsychological assessment is to determine differences in cognitive functioning in clinical settings. Carr et al. [50] used poset classification models of neuropsychological test data to classify samples to detailed cognitive profiles using ADNI2 and AIBL data. In the risk of disease progression of Alzheimer’s disease for individiuals with mild cognitive impairment (MCI), Tatsuoka et al. [171] suggest that poset-based modeling methods may be useful for providing more precise classification of cognitive subgroups among MCI for imaging and genetics studies, and for developing more efficient and focused cognitive test batteries. An order structure arises naturally with skill profiles. The flexibility to not necessarily assume that one state is greater than another is an appealing feature of posets[169]. In addition, posets are comprised of states, into which cases are classified, that are associated with distinct patterns of attribute strengths and weaknesses [120]. Posets have several advantages over conventional statistical methods for handling large numbers of polyfactorial neuropsychological test variables such as the ability to mimic the expert judgment of a clinical neuropsychologist for each case in a large sample. They are efficient since valid conclusions can be drawn based on relatively few measures and classification of large samples can be accomplished rapidly [119]. However, Jaeger et al. [119] noted the limitation in the poset model applied to large conventional neuropsychological test dataset is that it is restricted to a selected set of attributes, while additional important distinctions remain to be tested.
@@ 8.4 Other
The application of posets and lattice models is not limited to the afore- mentioned domains and tasks for data analysis. There have been explorations of applying posets to a variety of problems such as hypothesis management in large scale research projects [104]; adaptive testing for cognitive assessment [172]; machine Learning of reviewer’s preferences on papers [268], similarity learning use pose estimations [20], and many more data driven applications.
@@ 8.5 Some Selected Datasets
We classify the data sets into: Sustainable development dataset; Computer Vision Dataset; Natural Language
@@
  • In the study of pose estimation in [20] for the study of visual similarity learning using posets, three data sets were considered: The Olympic Sports dataset [331], The Leeds Sports Pose(LSP)[330], MPII Pose [329].
@@
Sustainability data have been subdivided in two datasets, the first comprising socio-economic indicators and the second comprising environmental data.
@@
  • A battery of biochemical, microbiological and bioassay tests were used to identify degraded or degrading sediments and waters [29, 31, 7]. The dat set is available in [82].
@@
8.6 Selected software algorithms and packages
@@
  • Python implementation of deep visual similarity learning using posets [336]. Based on [20].
  • [83] Given an edge-weighted poset, the algorithm in [336] constructs a total order such that nodes with smallest weights are adjacent. The algorithm works by attempting to order a set of words as closely as possible to their original surface realization in the UD corpus. Due to the fact that words may repeat in the sentence, each order is instead represented by a list of integers, and it is these lists of integers which are compared [336]. For example, assuming a target reference order of [1,2,3] for the red horse, the generated order of red the horse would be [2,1,3].
@@
  • The PARSEC is an An R package for poset-based evaluation of multidimen-sional poverty which was introduced in [11]. Its main goal is to provide socio-economic scholars with an integrated set of elementary functions for multidimensional poverty evaluation, based on ordinal information. The package is organized in four main parts
    1. 1.
      Data management.
    2. 2.
      Basic poset analysis.
    3. 3.
      Poset-based evaluation.
    4. 4.
      OPHI counting approach.
  • Poset algorithm for multidimensional data analysis [13].
  • Self Organizing Map(SOM)Self-organizing map algorithm[286], which is one of the most used tools for non-linear dimensionality reduction and pattern recognitio, in multidimensional data analysis. It has been used in:[12].
  • In [13] illustrate a non-aggregative algorithm new way of constructing sustainability synthetic indicators on multi-indicators systems.
  • A novel application of partial order theory for object rank correlation analysis of multiple variables is presented by the new software package PO Correlation [162] . The design is made transparent for rank correlation analysis by a detailed mapping of the rank relations between all objects.
  • The software WHASSE is applied for chemical monitoring data analysis [32, 34], medical data analysis [78].
@@
  • PyHasse Software [42, 51, 127] is for the pupose of the ordinal analysis of data matrices, identification and analysis of partial order relations as well as computing ranks.. It is used in [179] for the analysis of environmental data .
9 Future Research Directions
@@
Learning Algorithms A major challenge in data science is the identification of geometric structure in high-dimensional data. The structural understanding of data is very relevant for designing efficient algorithms for optimization and machine learning. Classically, the structure of data has been studied under Euclidean assumptions. The fundamental representation of any feature for any machine learning model is the vector and its multidimensional generalization which is the tensor. As such many machine learning algorithms and data analyis pipelines have been developed with the aim of computing vectors or matrices of real numbers. There are currently a wide range of well-studied tools and algorithms that assume such structure. However, many scientific fields study data with an underlying structure that can only be represented in non-Euclidean space. Such data can be found in graphs and many kinds of manifolds. There is currently growing research on developing machine learning models that work with data representations for features and embeddings, and or model parameters in the non Euclidean domains. Posets as a class of directed acyclic graphs have also begun to recieved attention with regards to developing appropriate learning algorithms. Recently, Wendler [180] introduced new methods for the fourier-sparse learning on data indexed by lattices and posets. The graph neural network class of algorithms also works on posets as specified in section 3. But there is not yet widespread use of graph neural networks relative to the number of its applications based on posets and lattices. The success of machine learning algorithms generally depends on data representation since as different representations encode different explanatory factors of variation behind the data [283]. There is enormous scope to develop new robust and efficient algorithms for learning the key structural characteristics of data represented as lattices and posets.
Posets and Topological Data Analysis Topological data analysis (TDA)[325] is a recent and fast-growing field in applied mathematics which provides a set of new topological and geometric tools to infer relevant features for complex data. There is a growing body of literature which connects posets to topology data analysis. This current article did not take into consideration the posets that appear in topological data analysis. There are many new applications which can arise from the study of toplogical analysis of partially ordered data.
@@
Learning to rank with partially ordered data Learning to rank refers to machine learning techniques for training the model in a ranking task. The prediction of structured outputs in general and rankings in particular has attracted considerable attention in machine learning in recent years, and different types of ranking problems have already been studied. There has been many ranking problems introduced in the context of data analysis of partially ordered data in sections 5 and 7. There is a great potential of introducing existing and new techniques of learning to rank for ranking problems with datasets from which can be derived a partial order structure such as big data for sustainable development etc. In particular, it can be useful to implement solutions from the various graph based[339, 338] approches of learning to rank for this purpose or developing new learning to rank algorithms for partially ordered data.
@@ 10 Conclusion
This article can be considered the first exhaustive review summarizing the applications of posets in relation to data analysis and machine learning. The vast use of poset theory in many domains of machine learning and data analysis demonstrates their relevance for a long time to come. We have examined the growing adoption of poset methodology in several areas of machine learning and deep learning. Posets theory was also systematically categorized into multidimensional, descriptive and explorative data analysis from many perspectives. The practical applications of posets was drawn from a broad range of domains. The main contributions of this work is to highlight the deep connections poset theory has with data analysis and machine learning. It is expected that this article will provide a basic and in-depth understanding of the applications of posets in data science and the discovery of new directions for future research on the subject.

### v2

@@ 8.3 Neurocognition Modelling
Finite partially ordered sets are natural models for cognition as it is reasonable to assume that some cognitive states have higher levels of functionality than others[272, 273, 271, 275, 276]. Additionally, finite partially ordered classification models are useful for many statistical applications including cognitive modelling. In particular, a data analytic framework for implementing latent finite partially ordered classification models introduced by Tatsuoka[272], provides useful methods for evaluating cognitive applications that are latent and complex. Poset models are flexible and can become quite rich and complex, enabling them to be effective models for describing response phenomena from educational test data or neuropsychological assessment data[272, 274, 279]. An objective of neuropsychological assessment is to determine differences in cognitive functioning in clinical settings. Carr et al.[92] used poset classification models of neuropsychological test data to classify samples into detailed cognitive profiles using ADNI2(Alzheimer’s Disease Neuroimaging Initiative) and AIBL(Australian Imaging, Biomarker & Lifestyle ) datasets. In the risk of disease progression of Alzheimer disease for individuals with mild cognitive impairment (MCI), Tatsuoka et al. [277] suggest that poset-based modeling methods may be useful in providing more precise classification of cognitive subgroups among MCI for imaging and genetics studies, and for developing more efficient and focused cognitive test batteries. An order structure arises naturally with skill profiles. The flexibility to not necessarily assume that one state is greater than another is an appealing feature of posets[275]. In addition, posets are comprised of states, into which cases are classified, that are associated with distinct patterns of attribute strengths and weaknesses [172]. Posets have several advantages over conventional statistical methods for handling large numbers of polyfactorial neuropsychological test variables such as the ability to mimic the expert judgment of a clinical neuropsychologist for each case in a large sample. Posets are efficient in these tasks since as valid conclusions can be drawn based on relatively few measures, and classification of large samples can be accomplished rapidly [171]. On the other hand, Jaeger et al.[171] noted that the primary limitation in poset modeliing of large conventional neuropsychological test datasets is based on its restriction to a selected set of attributes, while additional important distinctions remain to be tested.
@@ 8.4 Other
The application of posets and lattice theory are not limited to the aforementioned domains and tasks for data analysis. Posets have also been used in: 1) hypothesis management in large scale research projects [147]; 2) longitudinal data analysis[113]; 3) social macroeconomics analysis[93] ; 4) adaptive testing for cognitive assessment [278]; 5) psychometrics[166]; 6) medical statistics [261]; 7) machine learning of reviewer’s paper preferences [96]; 8) visual similarity learning using pose estimations [31]; 9) gender analysis [85]; 10) Pattern Mining [268] etc.
@@ 8.5 Some Selected Datasets
  1. 1.
@@
    • The Olympic Sports dataset [317], The Leeds Sports Pose(LSP)[316], MPII Pose [315]. They are employed in the study of visual similarity learning from pose estimation using posets.[31].
@@
    The sustainable datasets are categorized into socio-economic and environmetal datasets.
@@
    • Data on crime rates and their ranking for sixteen American cities is described in Levy[207].
  4. 4.
@@
    • A battery of biochemical, microbiological and bioassay tests were used to identify degraded or degrading sediments in waters [47, 50, 11]. The dataset is available in [115].
@@
    • Multiple indicator data for stream channel stability at bridge crossing is described in [227].
  5. 5.
    Other datasets:
    • (i)Data covering a time period up to the end of 2022 on the historical head-to-head matches of six professional tennis players.(ii) Data on educational testing from 15 OECD countries in.reading comprehension based on test performance in 2015 from the Programme for International Student Assessment (PISA). Both of these are described in [270] for the purpose of partial ranking of tennis players and total ranking of educational systems respectively.
    • A mass spectroscopy dataset consisting of 11 phosphoproteins and phospholipids containing approximately 854 measurements of abundance levels in an observational setting described in the supplementary material in [252]. It is used in Taeb et al. [270] for the purpose of learning causal relations and structures in proteins.
    • UCI machine learning repository [313], used in [217].
    • Poset/Hypergraph Currvature Datasets[324], used in [303] to perform an empirical study involving computation and analysis of the Forman–Ricci curvature of hyperedges in 12 real-world hypergraphs.
8.6 Selected Algorithms and Software Packages
  1. 1.
@@ 8.6 Selected Algorithms and Software Packages
    • Python implementation of deep visual similarity learning using posets [322]. Used in [31].
    • Given as input an edge-weighted poset, the algorithm in [116] constructs a total order such that nodes with smallest weights are adjacent. The algorithm works by attempting to order a set of words as closely as possible to their original surface realization in the Universal Dependencies(UD) corpus. Due to the fact that words may repeat in the sentence, each order is instead represented by a list of integers, and it is these lists of integers which are compared [116]. For example, assuming a target reference order of [1,2,3] for the red horse, the generated order of red the horse would be [2,1,3].
@@
    • A learning algorithm for discovering partial orders from sequences of events is described in [212].
    • The learning algorithm of pertinent concept described in [217] uses the Adaboost algorithm with formal concept analysis on a learning dataset to discover lattice concepts used for classification rules.
    • An algorithm framework for integrating base classifier into concept node of concept lattice is described in [301].
    • Greedy sequential algorithm for model selection described in [270]. It is used in partial ranking of tennis players, total ranking in educational systems, and causal structure learning on proteins dataset.
  2. 2.
@@
    The Hasse Diagram Technique (HDT) described in section 7 provides a huge collection of methods that are simple but tend to be complicated if the number of elements in the poset increases. To address this chalenge, special software packages have been developed to support the HDT. Some of these are described as follows.
    • The PARSEC[15, 122] is an R package for poset-based evaluation of socio-economic data. Its main goal is to provide socio-economic scholars with an integrated set of elementary functions for multidimensional poverty evaluation based on ordinal information. The package is organized in four main parts
      • i) Data management.. ii) Basic poset analysis. iii) Poset-based evaluation. iv) OPHI counting approach.
    • An algorithm based on posets for multidimensional data analysis is described in [17].
    • Self-organizing map algorithm(SOM)[314] is a popular tool for non-linear dimensionality reduction and pattern recognition. It is used for multidimensional data analysis described in [16].
    • A non-aggregative partial order algorithm for the construction of sustainability synthetic indicators on multi-indicators systems is described by Arcagni et al. [17].
    • An application of partial order theory for object rank correlation analysis of multiple variables is implemented by the software package PO Correlation [267] . The design is made transparent for rank correlation analysis by a detailed mapping of the rank relations between all objects.
    • The software WHASSE is applied for chemical monitoring data analysis [51, 53, 157], medical data analysis [111].
@@
    • PyHasse software [62, 66, 198] is used for the purpose of ordinal analysis in data matrices, identification and analysis of partial order relations as well as in computing ranks. Furthermore, it has been used in [293, 55, 290] for the analysis and evaluation of environmental data and in [87] for multi-criteria decision analyses.
    • ProRank[238] software for partial order ranking. It is used in the evaluation of environmental databases [292].
9 Future Directions
@@
Learning Algorithms A major challenge in data science is the identification of geometric structure in high-dimensional data. The structural understanding of data is very relevant for designing efficient algorithms for optimization and machine learning. Classically, the structure of data has been studied under Euclidean assumptions. The fundamental representation of the features for any machine learning model is the vector and its multidimensional generalization which is the tensor. As such many machine learning algorithms and data analyis pipelines have been developed with the aim of computing vectors or matrices of real numbers. There are currently a wide range of well-studied tools and algorithms that assume such structure. However, many scientific fields study data with an underlying structure that can only be represented in non-Euclidean space. Such data can be found in graphs and many kinds of manifolds. There is currently growing research on developing machine learning models that work with data representations for features and embeddings, and or model parameters in the non Euclidean domains. Posets as a class of directed acyclic graphs have also begun to recieve attention with regards to developing appropriate learning algorithms. Recently, Wendler [296] introduced new methods for the fourier-sparse learning on data indexed by lattices and posets. The graph neural network class of algorithms also works on posets as specified in section 3. But there is not yet widespread use of graph neural networks relative to the number of i applications based on posets and lattices. The success of machine learning algorithms generally depends on data representation since as different representations encode different explanatory factors of variation behind the data [33, 112]. There is enormous scope to develop new robust and efficient algorithms for learning the key structural characteristics of data represented as Hasse graphs of lattices and posets from a graph based perspective. There is also the need to develop specialized measures of model evaluation metrics for the use in any potential poset learning algorithms.
Posets and Topological Data Analysis Topological data analysis (TDA)[94] is a recent and fast-growing field in applied mathematics which provides a set of new topological and geometric tools to infer relevant features from complex data. In topological data analysis[178], the shape of a dataset is often encoded into a system of vector spaces and linear maps over a partially ordered set. There is a growing body of literature which connects posets to topological data analysis [177, 72, 40, 281]. The current survey did not take into consideration the posets that appear in topological data analysis.More recently, there has also been the introduction of posets in the construction of topological deep learning models[155]. There are many new applications which can arise from the topological analysis of partially ordered data and the development of their associated deep learning framework. There is the need to study situations of multiway interactions within data points endowed with a poset structure, by leveraging concepts such as Latent Topology Inference (LTI) [30].
@@
Learning to rank with partially ordered data Learning to rank refers to machine learning techniques for training the model in a ranking task. The prediction of structured outputs in general and rankings in particular has attracted considerable attention in machine learning in recent years, and different types of ranking problems have already been studied. Several ranking problems have been introduced in the context of data analysis of partially ordered data in sections 5 and 7. There is a enormous potential of introducing existing and new techniques of learning to rank, for ranking problems with datasets from which can be derived a partial order structure such as big data for sustainable development etc. In particular, it can be useful to implement solutions from the various graph based[142, 117] approaches of learning to rank for this purpose or developing new learning to rank algorithms for partially ordered data.
@@ 10 Conclusion
This article can be considered the first exhaustive review summarizing the applications of posets in relation to data analysis and machine learning. The vast use of poset theory in many domains of machine learning and data analysis demonstrates their relevance for a long time to come. We have examined the growing adoption of poset methodology in several areas of machine learning and deep learning. Poset theory was also systematically categorized into multidimensional, descriptive and explorative data analysis from many perspectives. The practical applications of posets was drawn from a broad range of domains. The main contributions of this work is to highlight the deep connections poset theory has with data analysis and machine learning. It is expected that this article will provide a basic and in-depth understanding of the applications of posets in data science and the discovery of new directions for future research on the subject.